OpenStudy (anonymous):

how in the world do i solve the inverse function of this f(x)=2+3^4x-5

OpenStudy (anonymous):

y=2+3^{4x-5} y-2=3^{4x-5} Ln(y-2)=Ln(3^{4x-5}) Ln(y-2)={4x-5}*Ln(3) [Ln(y-2)/Ln(3)]=4x-5 [Ln(y-2)/Ln(3)]+5=4x x=4*{[Ln(y-2)/Ln(3)]+5}